(3 parts for the answers)

Cystic Fibrosis Cystic fibrosis is an autosomal recessive disorder that affects 1 in 3 000 newborns with a Caucasian background. It causes thick mucus build up in various organs and can cause damages and problems in respiratory and digestive systems.

--What is the frequency of the cystic fibrosis allele in the Caucasian population? Show all your work and express your answer as a value between 0 and 1 rounded to two decimal places

--What percentage of the Caucasian population would be carriers for cystic fibrosis? Show all your work and express your answer rounded to two decimal places

--If two individuals are carriers of the cystic fibrosis allele, what is the probability that they would have a girl without cystic fibrosis? Show all your work (including the Punnett square) and express your answer as a value between 0 and 1 rounded to two decimal places.
Answer #1

According to Hardy-Weinberg equilibrium sum of all the allelic frequency of a gene is always 1 and sum of all the genotypic frequency of all the genotype is always 1.

So p+q =1

p2+ 2pq+q2 =1

here p = frequency of dominant allele

q is the frequency of recessive allele.

q2 = proportion of cystic fibrosis in the population

Given that q2 = 1/3000 = 0.00033

Hence q = 0.057

Also p+q =1

so p = 1-0.057

or p = 1-0.057

Hence p =0.943

Frequency of heterozygous or carrier will be 2pq

or 2*0.057*0.943 = 0.1075

So the percentage of population which are carrier for cyctic fibrosis will be = 0.1075*100 = 10.75%